[Hallc_running] [New Logentry] RC Daily Update

bduran at jlab.org bduran at jlab.org
Tue Feb 18 14:00:03 EST 2020

Logentry Text:
The accelerator was down from 8:30 am Monday morning to around 5 am on Tuesday for the Wien flip. A nice explanation about Wien flip can be found in this entry. https://logbooks.jlab.org/entry/3786542
During the downtime, we took an opportunistic controlled access when some target group people (Jian-Ping, Junhao and Brad) installed rotation stages and camera in the laser optics box and, mounted the longitudinal QWPs to the rotation stages (https://logbooks.jlab.org/entry/3786216 https://logbooks.jlab.org/entry/3786229). They also added fibers to the laser spectrometer and set them up in a way longitudinal and transverse spectrometer signal can be separated by disconnecting corresponding fibers (https://logbooks.jlab.org/entry/3786230). After the work in hall was completed around 9:30 pm Murchanna set longitudinal QWP angles to their optimized values (https://logbooks.jlab.org/entry/3786342). An NMR done at 1:50am gave 39.4% longitudinal polarization!!!(https://logbooks.jlab.org/entry/3786300). After the beam came back, the shift crew followed the beam checkout procedure. The first and second harp scans indicated pretty wide beam width (https://logbooks.jlab.org/entry/3786!
 379 https://logbooks.jlab.org/entry/3786382 https://logbooks.jlab.org/entry/3786399 https://logbooks.jlab.org/entry/3786400). After the operator did their magic, the third harp scans (https://logbooks.jlab.org/entry/3786410 https://logbooks.jlab.org/entry/3786411) gave reasonable beam size to allow us accepting CW beam to the hall. Raster was set to 4.0 mm and we asked for 30 uA to take production data with 3He with longitudinal spin direction. We're getting maximum 27 uA.

At MCC meeting, the crew chief pointed out that they are having issues to send the desired beam to hall A and C and want to do injector matching (5 MeV and injector to north linac) that wasn't done after the Wien flip to fix the beam problems. Starting from 8:30 am today, we were down until 12:40 pm when the beam was back we followed the short beam check out procedure and asked for harp scans (https://logbooks.jlab.org/entry/3786665 https://logbooks.jlab.org/entry/3786666) which indicated projected sigmas at the target: 0.4596/0.3148. Following the scans, we did an NMR --> 47% longitudinal polarization (https://logbooks.jlab.org/entry/3786686). Then we asked for 4.0 mm raster size and 30 uA and started taking data with 3He longitudinal. We will be sitting there for a while. Next planned Moeller measurement will be done tomorrow during the day shift.

We will have an RC meeting today at 4:15 pm at the second floor of the counting house. Here is how to connect remotely:
To join via Room System:
Video Conferencing System: bjn.vc -or-
Meeting ID : 861439920
To join via phone :
1) Dial:
+1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3786732
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200218/8af60894/attachment.html>

More information about the Hallc_running mailing list